OpenStudy (anonymous):

Find the following indefinite integral:

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ \tan \sqrt{x} }{ \sqrt{x} } dx\]

zepdrix (zepdrix):

\[\int\limits\limits_{}^{} \frac{ \tan \sqrt{x} }{ \sqrt{x} } dx \qquad = \qquad \int\limits\limits_{}^{} \tan \sqrt{x}\left(\frac{1}{\sqrt{x}} dx\right)\]See how I just took the tangent off of the top of the fraction? From here you can do a `U substitution`, let \(u=\sqrt x\)

OpenStudy (anonymous):

i was letting u=tanrootx

zepdrix (zepdrix):

yah that's no good! :) heh

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

so my du--1/x^2?

zepdrix (zepdrix):

\[du=\frac{1}{2\sqrt x}\]You might want to convert your square root to a rational expression if you don't have this derivative memorized. What I mean is, change it to x^1/2 so it's easier to differentiate.

zepdrix (zepdrix):

woops I didn't write the dx on the right* it should be there.

OpenStudy (anonymous):

okay so thats my du.

zepdrix (zepdrix):

\[\large u=x^{1/2} \qquad \rightarrow \qquad du=\frac{1}{2}x^{-1/2}dx \qquad \rightarrow \qquad du=\frac{1}{2\sqrt x}\]Yup, there is our du. See how the right side ALMOST matches the parthenthesis'd part in our integral?? Is there a way we can make them match, so we can apply the substitution?

zepdrix (zepdrix):

I dropped the dx again!!! ahhhh

OpenStudy (anonymous):

can we take out 1/2?

zepdrix (zepdrix):

\[\large du=\frac{1}{2}\color{orangered}{\frac{1}{\sqrt x}dx}\]And here is our integral again,\[\int\limits\limits\limits_{}^{} \tan \sqrt{x}\left(\color{orangered}{\frac{1}{\sqrt{x}} dx}\right)\]

zepdrix (zepdrix):

Take out? like on a date?? jk, yah i think you have the right idea maybe :) you want to multiply both sides by 2, to get rid of that 1/2 on the right.

zepdrix (zepdrix):

Ahhh sorry I have to go to school :C If you're still stuck on this, i'll be @sirm3d or someone can help you finish it!! c:

OpenStudy (anonymous):

i meant taking out a 1/2 outside the inegral sign haha.

OpenStudy (anonymous):

thanks for your help!

OpenStudy (sirm3d):

after \[u=\sqrt x\] square both sides \[u^2=x\\2udu=dx\\\int \frac{\tan \sqrt x}{\sqrt x}dx=\int \frac{\tan u}{u}(2u\mathrm du)\]